Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 309     Entry time: 17 Aug 2023, 11:08
Author: Teguh Citra Asmara 
Group: SCS 
Subject: Afternoon shift summary on 16.08.2023 

Achievements:

  1. tr-RIXS at 4 mJ/cm2 for energies A, B, C, and D (pre-edge)
  2. Checked and optimized FEL alignment, t0 and overlap with PP laser after reset of machine synchronization
  3. Long delay scans for Peak A and C, up to 100 ps
    Problems:

  4. Need to reset machine synchronization by BKR

  5. Camera scene got stuck, was resolved by DOC

ELOG V3.1.4-7c3fd00